40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog, Page 16 of 18  Not logged in ELOG logo
ID Date Author Type Category Subject
  131   Wed Sep 28 10:26:58 2016 ranaGeneralGeneralreferences on ideal glasses

Some easy to read reviews on thin films and ideal glass for people getting started in this Q measuring game.

  1. M. D. Ediger, in PNAS  (2014), pp. 11232–11233.

  2. A. J. Leggett and D. C. Vural, arXiv cond-mat.dis-nn, arXiv:1310.3387 (2013).

  3. L. Berthier and M. D. Ediger, arXiv cond-mat.mtrl-sci, 40 (2015). Phys. Today.

  4. G. Parisi and F. Sciortino, Nature Materials 12, 94 (2013).

  5. S. Singh, M. D. Ediger, and J. J. de Pablo, Nature Materials 12, 139 (2013).

  130   Tue Sep 27 15:10:59 2016 GabrieleGeneralMeasurementsRing-down of etched disk

Quiet (roughing pump off, lights off): 60 seconds from
PDT: 2016-09-27 15:05:30.805667 PDT
UTC: 2016-09-27 22:05:30.805667 UTC
GPS: 1159049147.805667
Follows excitation and ring-down with QPD autocentering (10 seconds interval). Centering is good starting 215 seconds after the time above.

There is a drift in X, corrected by the picomotor.

The spectrum of both QPD normalized signals looks quite bad. Maybe there's some scattered light issue.

  129   Tue Sep 27 14:54:53 2016 GabrieleElectronicsConfigurationEpics values now saved to frames

Apparently, there was a mismatch in the configuration, and DAQD was adding a wonderful 16 Hz comb all over the spectrum.

I stopped the processes, but couldn't restart x3cr1. It turned out that I can't save a channel to frames with a sampling frequency lower than 256 Hz. I changed the model, recompiled and restarted. Now the 16 Hz is gone.

Quote:

Now EPICS values are saved to frames, but they are all zero! I noticed that we always had the same problems with the cymac2 too.

So for the moment being I set up daqd to save X_NORM_IN1 and Y_NORM_IN1 at 32 Hz. In this way I can monitor the QPD centering.

Quote:

I was looking at some past trend data and discovered that EPICS values were not written to the frames. I added the following two lines to /opt/rtcds/tst/x3/target/fb/master to fix this:

/opt/rtcds/tst/x3/chans/daq/X3EDCU_CR1.ini
/opt/rtcds/tst/x3/chans/daq/X3EDCU_TST.ini

 

 

  128   Tue Sep 27 08:53:46 2016 GabrieleGeneralVacuumEtched disk installed

Installed the etched disk: using manually the centering ring allowed me to get the beam on the QPD. A couple of taps to the disk were enough to get the beam centered.

Pump down started at 8:52am

  127   Mon Sep 26 15:59:09 2016 GabrieleOpticsDaily ProgressQPD auto centering

We have a few motorized mounts (with New Focus picomotors) and one controller (an old New Focus 8753, six axis total) that I connected with a makeshift null modem cable to the laboratory workstation (better cabling and power supply coming soon).

I wrote a couple of python scripts that can be used to continuosly read out the QPD values and move the picomotors if needed. It's wortking quite well, so we should be able to use it in the future to keep the QPD centered during the measurement. 

The scripts are in the ~/CRIME directory. Launch the function center() in the autocenter.py script.

  126   Mon Sep 26 15:54:01 2016 GabrieleFacilityConfigurationMoved vacuum controllers

I moved the turbo pump controller out of the clean room. Also, I installed the gauge controller on the Cymac rack.

  125   Fri Sep 23 08:33:39 2016 GabrieleElectronicsConfigurationEpics values now saved to frames

Now EPICS values are saved to frames, but they are all zero! I noticed that we always had the same problems with the cymac2 too.

So for the moment being I set up daqd to save X_NORM_IN1 and Y_NORM_IN1 at 32 Hz. In this way I can monitor the QPD centering.

Quote:

I was looking at some past trend data and discovered that EPICS values were not written to the frames. I added the following two lines to /opt/rtcds/tst/x3/target/fb/master to fix this:

/opt/rtcds/tst/x3/chans/daq/X3EDCU_CR1.ini
/opt/rtcds/tst/x3/chans/daq/X3EDCU_TST.ini

 

  124   Fri Sep 23 08:11:48 2016 GabrieleMechanicsDaily ProgressTest of the disk retaining ring motion

In brief, it doesn't work. The magnets and coils are strong enough to push up the ring with a sample inside, but the friction with the three alignment pins is too large and random, so when the current to the coils is increased slowly, the ring doesn't move up smoothly (see first attached video). On the other hand, if the current is switched on abruptly, the ring shoot to the top and stays there. However, if a disk is placed on the support, it is ejected out (see second video). When the current is cut (smoothly or abruptly) the ring doesn't alway comes back to the bottom, but sometimes it stays stuck inclinded.

On the positive side, we probably don't need such a complicated system:

  1. in all the pump down I've done so far (ten or more), the disk never moved
  2. the ring is very useful, even when used manually, to find the initial centering of the disk: if we machine three small aluminum wedges that can be put under the ring to keep it raised (or three set screws), it can be used to place down the sample in a roughly centered position, that has always been good enough to get the beam almost back into the QPD.

Links to the two videos:
Video1 Video2

  123   Thu Sep 22 15:00:49 2016 GabrieleElectronicsConfigurationEpics values now saved to frames

I was looking at some past trend data and discovered that EPICS values were not written to the frames. I added the following two lines to /opt/rtcds/tst/x3/target/fb/master to fix this:

/opt/rtcds/tst/x3/chans/daq/X3EDCU_CR1.ini
/opt/rtcds/tst/x3/chans/daq/X3EDCU_TST.ini
  122   Thu Sep 22 13:39:40 2016 GabrieleOpticsConfigurationUpdated design for the new C.Ri.Me. setup

Today I measured the amount of space available on the table for the new (4-fold) C.Ri.Me. setup. It's 1050 x 1220 mm, with the table hole in it. 

So I updated the optical layout to fit into this space, and optimized the telescope to have a beam spot on the QPD of the order of 350 um. The average lever arm length is 1.5 m, so the optical gain will be about 7000 /rad.

 

Attachment 1: crime_v2.pdf
crime_v2.pdf
  121   Thu Sep 22 09:49:21 2016 GabrieleGeneralGeneralRing-down of etched disk

QPD centerd, quiet data, light off, one minute from

PDT: 2016-09-22 09:46:29.393609 PDT
UTC: 2016-09-22 16:46:29.393609 UTC
GPS: 1158598006.393609

Excitation (2kV) stopped at

PDT: 2016-09-22 09:49:03.784165 PDT
UTC: 2016-09-22 16:49:03.784165 UTC
GPS: 1158598160.784165

 

  120   Wed Sep 21 14:42:47 2016 GabrieleGeneralGeneralSerialized (etched) disk installed

Just for fun, I installed the disk thas has been etched in the center with "1234". I figured out that the ESD PCB was probably too close to the disk, so I moved it a bit up.

Pump down started at about 2:38pm LT.

  119   Wed Sep 21 08:12:13 2016 GabrieleGeneralGeneralExcitation and ring down

New excitation (2000V) at about 8:06am. Had to recenter the QPD again after the excitation.

Engaged the 500Hz high pass filter on the ESD filter bank. New excitation ended at 8:11am. Amplitude 1000 V. Recentered the QPD at 8:11:35am

  118   Tue Sep 20 16:43:58 2016 GabrieleGeneralGeneralExcitation and ring down

The pressure is at abour 3e-6 Torr. I centered the QPD and started an excitation. The HV amplifier manual states that the driver can source both positive and negative voltage, so this time I didn't add any offset, but simply drove with 1000 V peak to peak. After the excitation the QPD was slightly miscentered in X and I had to manually recenter it.

Good data starting from

PDT: 2016-09-20 16:38:09.330642 PDT
UTC: 2016-09-20 23:38:09.330642 UTC
GPS: 1158449906.330642

NOTE: it's a good idea to take a look at both the X and Y signals for each mode. Some of them look stronger in Y than in X. So far I only used X.

  117   Tue Sep 20 16:12:09 2016 GabrieleGeneralGeneralPump down

As a test, I installed MO1 (the disk with the burn mark, used for the first edge laser polishing test) and started pumping down. Roughing pump on at 3:05pm, turbo pump on at 3:16pm.

  116   Tue Sep 20 15:34:00 2016 GabrieleOpticsDaily ProgressImproved optical lever layout

Goal

Improve the optical setup, by increasing the response of the QPD to disk motion.

The old configuration

In all my previous measurement the optical lever was as simple as possible: no lenses were used, and therefore the beam was free to expand over all its path. The estimated arm lever from the disk to the QPD was 1030 mm.

QPD response to disk angular motion

The response of the QPD can be characterized with its optical gain in 1/rad, which is how much the normalized signal (difference / sum) changes for one radians of motion of the disk. This is the product of two parts:

  1. the gain from angular motion of the disk to beam spot motion on the QPD. In the simple case of free propagation this is 2L, where L is the distance from the disk to the QPD, and the factor 2 is due to the fact that the beam deflection is the double of the disk angular motion. If there is a telescope in between the disk and the QPD, it is easy to compute the total ray transfer matrix:
    \begin{pmatrix} x_{QPD}\\ \theta_{QPD} \end{pmatrix}= \begin{bmatrix} A & B\\ C & D \end{bmatrix} \begin{pmatrix} 0\\ 2\theta_{disk} \end{pmatrix}
    Then the gain is simply the B element of the matrix.
  2. the response of the QPD normalized signal to beam motion. This depends only on the beam spot radius w on the QPD. It can be computed by simple gaussian integration, and in the approximation of small beam motion, it is given by the following expression:
    g = \frac{2}{w}\sqrt{\frac{2}{\pi}}

In the case of the old configuration, the beam spot size on the QPD was measured to be about 1.5 mm in radius, so the optical gain is of the order of 1900 /rad.

Laser beam profiling

Since I wanted to improve the optical setup, I first needed to measure the beam coming out of the HeNe laser. I used the WinCam beam profile and a Newport rail to measure the beam X and Y sizes at different positions.

The measurements are not the best ever, but I can still get a fit for the evolution of the gaussian beam, as shown in the plot below. The beam waist is 254 um, located 340 mm behind the laser output (inside the laser tube).

Design of the improved setup

I decided to try a brute force algorithmic optimization for the optical gain. I allow two lenses between the laser and the disk and two lenses between the disk and the QPD. I wrote a MATLAB script that picks the four lenses from a list of all those available (I have a Thorlabs LSB02-A lens kit). For each combination of lenses, MATLAB moves them around into pre-defined ranges, and try to find the maximum value of the QPD total optical gain, which is the product of the factor g above and of the B element of the ray tracing matrix.

It turned out that the best optical gains could almost always be obtained by making the beam huge on the disk (5-10 mm radius) and tiny on the QPD (tens of microns). This is not a good solution. So I decided that the beam on the disk must be smaller than 2mm in radius and the beam on the QPD must be larger than 200 microns. I enforced those limits into the optimization code by weighting the gain with a function which is one in the allowed range, and then quickly drops to zero when either of the beam sizes fall out of the allowed range.

The script ran for about half hour and gave me a lot of possible options. After some inspections, I decided to use the following one, which uses only one lens between laser and disk, and two between the disk and the QPD. Distances and focal lengths are shown below. Note that the first distance (laser to first lens) is from the laser beam waist to the lens, so the actual distance must take into account that the waist is estimated to be 340mm into the laser.

With this configuration the optical gain is computed to be 17000 /rad, or about 9 times larger than the original setup. The beam radius on the disk is 1 mm and on the QPD is 0.23 mm.

Implementation

First of all I measured some distances:

  • from the inner side of the viewport to the disk: 420 mm
  • viewport thickness: 12 mm, which is about 18 mm optical length considering n~1.5
  • so from the input to the chamber to the disk: 438 mm
  • from the viewport to the upper external periscope mirror center: 110 mm
  • distance between the periscope mirror centers: 275 mm

Using these distanced I build the designed optical setup. Some remarks on the procedure

  • I first aligned the laser beam to be horizontal, then added the first lens and centered it by ensuring no beam shift far away from the lens
  • I first aligned the periscope to get the beam roughly centered on the inner 45 degrees mirror, and then roughly centered on the black glass
  • Then I put a small container with water inside the chamber, on top of the black glass. I aligned the inner mirror and the periscope so that the beam coming back from the horizontal water surface was perfectly overlapped with the input beam. I used an iris on the input beam path
  • Then I removed the water container and installed a test disk. I moved the disk around until I got the same beam position in output. This tells me that the disk is horizontal
  • Finally I moved the upper periscope mirror to separate horizontally the beam coming back, at the level of the table. The separation is large enough to allow me to pick up the outgoing beam with a mirror.

Here's a picture of the setup, with the optical path highlighted. 

 

  115   Fri Sep 16 14:06:43 2016 GabrieleGeneralGeneralPump had no effect on ring down measurements

Same plot as below, but this time with estimated 95% confidence intervals for the Q values, as obtained from the fit only.

Quote:

The plot below shows three measurements of the Q of the same disk: during the first two the roughing pump was on, while during the third it was off. No significant difference is visible in the Q values.

 

 

  114   Thu Sep 15 15:15:10 2016 GabrieleGeneralGeneralEdge effect on losses

I made a COMSOL model that can compute the distribution of elastic energy for each mode, dividing it into:

  • bulk and shear energies (integrated over the entire volume)
  • edge energy (integrated over the edge surface only)
  • surface energy (integrated over the top and bottom surfaces)

Then I used the measured Q values for the MO_101 disk and tried to see if I could reproduce it with the energy distribution. The first plot here shows that the loss angle of the disk (inverse of the Q) has a trend that is already quite well reproduced by the ratio on edge energy over total energy:

In particular the edge energy distribution is enough to explain the splitting of the modes in families. This fit is obtained assuming that the edge losses are uniform along the entire edge, and frequency independent. If we assume a "thickness" of the edge of the order of 1 micron, the loss angle is about 3.5e-3, which seems resonable to me since the edge is not polished.

Then I tried to improve the fit by adding also bulk, shear and surface losses. It turns out that shear is not very important, while bulk and surface are almost degenerate. The following plot shows a fit using only edge and surface losses:

 

The result is improved, expecially for the modes with lower loss angle. Again, assuming a surface thickness of 1 micron, the main surfaces have a loss angle of 1.3e-5, while the edge is 2.3e-3.

Including all possible losses gives a fit which is basically as good as the one above:

However, the parameters I got are a bit differentL: the surface losses are reduced to zero, while bulk dominates with a loss angle of 1.4e-4, and shear is not relevant. 

In conclusion, I think the only clear message is that the Q of our disks are indeed limited by the edge. The remaining differences are difficult to ascribe to a paritcular source. Since th disks are thin, I tend to ascribe them to the surface, which would imply that we are far from being able to see the bulk/shear losses. If I use only edge and surface losses, I found as expected that the polished main surfaces have much lower loss angle by a factor 200 or so.

 

  113   Thu Sep 15 11:23:13 2016 GabrieleFacilityGeneralChamber open

Turbo pump off and spinning down at 9:37am LT. Pumo completely stopped at 11:15am LT

Openend the chamber and removed the sample at ~11:20am LT

  112   Thu Sep 15 08:27:57 2016 GabrieleGeneralGeneralNew band-limited excitation

Same as in elog #110, but now the amplitude is proportional to frequency squared:

ampl = (x/x[0])**2
xx = multi_band_noise(bands, ampl, T=20, fs=65536)
n = AWGNoiseStream(4e-4*xx, channel='X3:CR1-ESD_EXC', rate=65536)
n.start()

Noise stopped at 8:27:40am LT.

Turbo pump off and spinning down at 9:37am LT. Pumo completely stopped at 11:15am LT

Openend the chamber and removed the sample at ~11:20am LT

  111   Thu Sep 15 08:13:20 2016 GabrieleGeneralGeneralPump had no effect on ring down measurements

The plot below shows three measurements of the Q of the same disk: during the first two the roughing pump was on, while during the third it was off. No significant difference is visible in the Q values.

  110   Wed Sep 14 20:17:33 2016 GabrieleGeneralGeneralBand-limited excitation

Excitation started at 20:15:30LT, 20 seconds long. The excitation is band-limited (10 Hz) centered around each of the predicted mode frequencies. Amplitude inversely proportional to the mode frequency.  The system was quiet before the excitation for many minutes.

For reference, here's the code used for the excitation:

from noise import *
from numpy import *
x = loadtxt('predicted_modes.txt')
bands = map(lambda x: [x-5,x+5], x)
ampl = x/x[0]
xx = multi_band_noise(bands, ampl, T=20, fs=65536)

n = AWGNoiseStream(1e-2*xx, channel='X3:CR1-ESD_EXC', rate=65536)
n.start()

 

  109   Wed Sep 14 13:08:05 2016 GabrieleGeneralGeneralRingdown with roughing pump off

At 1:05pm LT I stopped the roughing pump and started a ring-down measurement. Pump restarted at 2:18pm LT.

  108   Tue Sep 13 10:33:35 2016 GabrieleGeneralGeneralDisk excited

Disk excited with white uniform noise, amplitude 5 V, for some tens of seconds.

Excitation off at

PDT: 2016-09-13 10:32:23.887615 PDT
UTC: 2016-09-13 17:32:23.887615 UTC
GPS: 1157823160.887615

 

  107   Mon Sep 12 16:00:06 2016 GabrieleGeneralDaily ProgressFirst fully automated measurements of ring down

I finished the first version of the automation software to measure the ring down of the disk modes. I tested it with the new substrate that was installed yesterday. Here are some screenshots and a brief explanation of how it works.

It is based on a Python/Tk GUI, that can be launched on the workstation with the command ~/CRIME/crime.py

The main screen is similar to the following. Once a baseline spectrum is acquired, it is shown in the main panel:

The user should specify the folder and prefix of the result files, and other parameters related to the excitation. The when the "Excite and ring down..." button is pressed, here's what happens

  1. If a baseline spectrum (before excitation) is not available, one is acquired with the specified parameters
  2. A broadband white excitation is applied with the selected amplitude and duration

  1. Another spectrum is taken. This is then whitened by dividing it with the baseline. This could be used directly to select the modes that have been excited. However, some parts of the noise floor are non stationary, so a second whitening is performed: the noise background is estimated by removing all lines, and it is then again divided out from the spectrum.
  2. All lines above a SNR threshold are then selected and shown in the main window together with the whitened spectrum: 

At this point the amplitude of the peaks are continuosly monitored (every second) and thei amplitude shown in a new window. The user can select a subset of the modes for the plotting.

There are some wandering peaks in the spectrum, so some of the peaks aren't actually modes that get excited. This is easily fixed in the post processing of the results. 

All peak amplitudes are saved to files in real time, so if you stop the GUI you'll have some partial results.

  106   Sun Sep 11 10:05:33 2016 GabrieleGeneralDaily ProgressNew substrate in the chamber

I installed one of the new substrates (with flats) into the chamber, and started the pumpdown at about 9:45am LT.

Before that, I removed the retaining ring: tomorrow I'm going to glue the magnet to it.

  105   Thu Sep 1 02:13:33 2016 GabrieleOpticsCharacterizationExpected frequencies for the 75 mm disks with flats

Here are the nominal parameters of the disk with flats

Parameter Value
Diameter (nominal) 75.0 mm
Thickness (nominal) 1.00 mm
Distance of flats from center of disk
(as measured by MO)
36.05 mm
Young's modulus (from G1601850) 72.3 GPa
Poisson's ratio (from G1601850) 0.164
Density 2202 kg/m^3

A COMSOL simulation gives the frequencies and mode shapes shown in the attached PDF file. Following the list of frequencies and a classification of the mode family (numer of radial nodes, number of azimuthal nodes in a half turn):

Frequency [Hz] Radial Azimuthal
1089.7 0 2
1109.1 0 2
1661.1 1 0
2501.3 0 3
2542.3 0 3
3837.7 1 1
3909.8 1 1
4359.9 0 4
4421.2 0 4
6656 0 5
6665.3 1 2
6732.7 0 5
6780 1 2
7283.9 2 0
9381.5 0 6
9470.6 0 6
10054.1 1 3
10210.2 1 3
11244.2 2 1
11414.2 2 1
12528.4 0 7
12655.4 0 7
13970.1 1 4
14148.9 1 4
15854.8 2 2
15856.8 2 2
16101.1 0 8
16227.8 3 0
16853.9 3 0
18387.2 1 5
18562.6 1 5
19984.9 0 9
20045.8 0 9
21067.2 2 3
21358.1 2 3
22300.9 3 1
22848.8 3 1
23270.5 1 6
23419.4 1 6
24374.1 0 10
24425.2 0 10
26793.8 2 4
27083.9 2 4
28533.5 1 7
28637.8 1 7
28692 3 2
28998.5 4 0
29228.3 0 11
29271.9 0 11
29870.1 4 0
33040 2 5
33301.4 2 5
33949.1 0 12
33986.5 0 12
Attachment 1: allmodes_aggregated.pdf
allmodes_aggregated.pdf allmodes_aggregated.pdf allmodes_aggregated.pdf allmodes_aggregated.pdf allmodes_aggregated.pdf allmodes_aggregated.pdf
  104   Tue Aug 23 09:44:11 2016 GabrieleGeneralGeneralMeasurements on MO-02

Two good ring-downs measurements were performed on MO-02. The first one was already reported in a previous elog entry. I performed another measurement, and refined the mode identification. I think I had misidentified some modes in my previous analysis. The following plot shows the difference between the modes as predicted by COMSOL and as measured. A clean quadratic trend is visible and fitted:

Here's the spectrum with all the modes:

And the updated Q measurement plot:

A second ring down was measured on Monday morning . Here are the relevant plots:

This is the same disk as before, but almost all Q values are systematically higher. Here's a direct comparison:

I'm not sure what changed between the two measurements, except for a re-alignment of the QPD. The disk might have moved a bit...

Attachment 4: spectrum_with_modes_12_01pm.png
spectrum_with_modes_12_01pm.png
  103   Tue Aug 23 09:21:25 2016 GabrieleGeneralGeneralStatus of Mark Optics disks

We initially received 20 disks (75 mm diameter, 1 mm thickness) from Mark Optics. Here's their status as of today

  • MO-01 was annelaed and measured at LMA. It was then taken back to Mark Optics by Julie Houser at the time of her visit
  • MO-02 was initially used to test the "dirty" measurement apparatus, then it was installed into the new chamber and measured on Saturday August 20th
    Then this disk was used as a guinea pig for the laser edge polishing, and it's now in a pretty bad shape
  • MO-03 was installed into the new chamber and measured a couple of times on Sunday and Monday. Then we laser polished the edges, with good results, although there are small damages on the surface. It has been installed back into the chamber and a couple of rind-down measurement carried out on Monday and Tuesday.
  • MO-04 is still untouched

All the other disks have been sent back to Mark Optics to grind out flats. 

  102   Tue Aug 23 08:46:04 2016 GabrieleGeneralDaily ProgressDaily tests

MO03 - edge polished:

Turbo off, QPD centered, before excitation (60 seconds)

PDT: 2016-08-23 08:42:54.514987
PDT UTC: 2016-08-23 15:42:54.514987
UTC GPS: 1156002191.514987

Excitation (white uniform noise, amplitude 5 V)

PDT: 2016-08-23 08:45:01.007626 PDT
UTC: 2016-08-23 15:45:01.007626 UTC
GPS: 1156002318.007626

Clean data for ring-down

PDT: 2016-08-23 08:45:46.448949 PDT
UTC: 2016-08-23 15:45:46.448949 UTC
GPS: 1156002363.448949

Restarted roughing pump, QPD got misaligned

PDT: 2016-08-23 10:00:29.259345 PDT
UTC: 2016-08-23 17:00:29.259345 UTC
GPS: 1156006846.259345

MO03 - edge polished:

Band-limited noise, +-10Hz around eahc nominal frequency, amplitude scaled based on the inverse of the peak height obtained with white noise. See attached code and plot

from numpy import *
from noise import *


x = loadtxt('/home/controls/Measurements/2016_08_23/mo_02_laserpolished_frequencies.txt')

freqs = x[:,0]
ampl = x[:,1]

bw = 10


bands = map(lambda x: [x - bw, x + bw], freqs)
a = 1 / (ampl/max(ampl))
a[a>50] = 50.


x = multi_band_noise(bands, a, 10, fs=65536)
x = x / 30

Ring down after:

PDT: 2016-08-23 11:07:02.661145 PDT
UTC: 2016-08-23 18:07:02.661145 UTC
GPS: 1156010839.661145

 

  101   Mon Aug 22 16:59:13 2016 Gabriele, AlastairOpticsDaily ProgressLaser polishing of the disk edges

We set up a test facility for laser polishing the disk edges, using the CO2 laser in the TCS laboratory. We focused the beam with a 10" focal length lens, and installed the disk on a "rotation stage" that we motorized with a hand drill. We used a HeNe optical lever and a small container with water to define the horizontal plane and adjusted the disk as well as we could.

We first tested the procedure on the MO02 disk, which is the one already scared with the electrostatic drive burn mark. This disk is now definitely in bad shape. However, we felt confident in our procedure, so we took out the MO03 disk that was into the measurement system and proceeded to laser polish the edges. Things went quite smothly. Unfortunately we added some small damages to the disk surface in a couple of spots where the CO2 laser went out of alignment and melted the fused silica support of the disk. The edge however looks quite good now.

Q measurement is on-going at the timw of writing

  100   Mon Aug 22 12:01:52 2016 GabrieleGeneralGeneralTests of the day

Disk excited at 12:01pm. Exited the room at 12:03pm.

Opened the chamber at about 2:30pm, got the disk out for edge polishing, installed it back at 3:30pm, pumping down at 3:40pm.

Stopped the roughing pump at 4:44:00pm (+60 seconds clean data, GPS 1155944657). Switched on the HV amplifier, excitation at 4:47:30pm. Recentered QPD, clean data from 4:48:30pm (GPS 1155944927)

After a first look at the data, it seems that something went wrong. I restearted the roughing pump and will pump overnight.  I found the QPD miscentered, so I centered it again.

Excited again at about 5:46:35pm. Clean data from 1155948460

  99   Mon Aug 22 08:29:24 2016 GabrieleGeneralDaily ProgressNew test

Restarted roughing and turbo pump at about 8:10am.

Quote:

I checked the status at about 5:20pm, the turbo pump was in error and spinning down, since the roughing pump has been off for about 1.5 hours.

I let the pump switch off.

Quote:

Roughing pump stopped at about 3:30:30pm. HV amplifier on at 3:33:30pm, excitation at 3:35:30pm. Recentered QPD at 3:36pm

 

 

  98   Sun Aug 21 17:41:34 2016 GabrieleGeneralDaily ProgressNew test

I checked the status at about 5:20pm, the turbo pump was in error and spinning down, since the roughing pump has been off for about 1.5 hours.

I let the pump switch off.

Quote:

Roughing pump stopped at about 3:30:30pm. HV amplifier on at 3:33:30pm, excitation at 3:35:30pm. Recentered QPD at 3:36pm

 

  97   Sun Aug 21 17:16:03 2016 GabrieleGeneralDaily ProgressRing down measurement with new disk

After fixing the setup, the measurement with the new disk looks great. After more than one hour the first two modes are still ringing down, meaning that the Q's are larger than 10 millions.

Here's the comparison of the spectrum before and after the excitation, with the identified modes:

I used about 4200 s of data to fit the ring downs. Most fits are good. In a couple of cases the peak splitting is large and the algorithm fails to fit the beats:

In summary, here are the Q values for all modes. Despite not being annealed, this disk shows very large Q's

  96   Sun Aug 21 15:36:16 2016 GabrieleGeneralDaily ProgressNew test

Roughing pump stopped at about 3:30:30pm. HV amplifier on at 3:33:30pm, excitation at 3:35:30pm. Recentered QPD at 3:36pm

  95   Sun Aug 21 13:45:56 2016 GabrieleOpticsDaily ProgressWierd behavior of ring downs and Improved setup

The last two ring downs I measured today showed a weird behavior of the lowest modes:

Although I'm not 100% sure, I suspect this is related to the fact that the beam reflected from the black glass was so close to the beam reflected by the disk that I could see interference.

So I broke vacuum and improved the setup, adding a peek washer below one edge of the black glass, to wedge it. In this way the reflection from the black glass is largely separated: it misses the upper periscope mirror and it is dumped on a black panel (together with the viewport reflection).

I realigned everything, installed back the disk and started pumping down at 1:30pm.

  94   Sun Aug 21 08:51:22 2016 GabrieleGeneralDaily ProgressNew disk installed

Vented the chamber. Installed a new disk (MO 03). The one I measured yesterday is now named MO 02 and it is the one with the "burnt mark" from the previous experiment (due to the electrostatic drive).

Startep roughing pump at 8:50am. Started turbo pump at 9.00am.

Excited the disk at 9:50:30am with white noise, amplitude 10 V. Pumps are still running, pressure is about 2e-6 Torr

At 11:10am I stopped the roughing pump, pressure is 1.4e-6 Torr. Exciting again the disk at 11:10:45am. At 12:25am I checked again the situation, since both measurements look quite weird, especially for the first couple of modes.

I think the reflection from the balck glass is interfering with the reflection from the disk. Probably I wasn't careful enough when I aligned the disk. At about 12:35 I stopped the turbo pump. I'm going to open the chamber and realign everything again.

  93   Sat Aug 20 22:18:26 2016 GabrieleGeneralDaily ProgressFirst measured Qs

Using the first ring down of the day (GPS 1155754513 + 3600 seconds), I computed the amplitude of each of the modes already identified, using a short FFT spectrogram (each FFT is 1 second long, overlap of 0.5 s).

Then I used the same code I developed at LMA to fit the ring down, including the beat between the unresolved mode pairs. The fit is versy sensitive to the initial conditions, so I had to fine tune them for each of the 20 modes. Still, all fits were successful with 30 minutes of work.

Here are all the fits:

And in summary all the measured Qs, which turned out to be larger than what I was expecting, considering that the disk is not annealed.

The analysis code in MATLAB is attached.

Attachment 22: first_ringdown.m
%% Compare data right before and right after the excitation
gps0 = 1155754000;
gps1 = 1155754513;
dt = 180;

c = nds2.connection('cymac3.ligo.caltech.edu', 8088);
data0 = c.fetch(gps0, gps0+dt, {'X3:CR1-X_NORM_OUT_DQ', 'X3:CR1-Y_NORM_OUT_DQ', 'X3:CR1-ESD_OUT_DQ'});
data1 = c.fetch(gps1, gps1+dt, {'X3:CR1-X_NORM_OUT_DQ', 'X3:CR1-Y_NORM_OUT_DQ', 'X3:CR1-ESD_OUT_DQ'});

x0 = data0(1).getData();
... 606 more lines ...
  92   Sat Aug 20 17:01:10 2016 GabrieleGeneralDaily ProgressFirst ring down measurements

Everything is working pretty well. This morning the pressure was about 1.2e-6 Torr. I connected the high voltage amplifier and I could drive the disk without problems.

I measured the beam shape and size at the QPD. We have about 50 uW, we see a TEM01-like mode due to the interference of the two disk surfaces (this is normal). The beam is about 3 mm in diameter. using this information and the estimated optical lever length of 1.2 m, I calbrated the QPD NORM signals in units of angular motion of the disk surface. The computation posted in CRIME_Lab/60 is actually wrong. I'll post the correct one later.

Injecting broadband white noise I could excite all the modes that are visible up to about 30kHz. I tuned the COMSOL model, by changing the thickess of the disk to 1.017 mm, to fit the frequency of the first few modes. Here are the modes I could measure:

Nominal frequency [Hz] Actual frequency [Hz]
1108 1109
2543 2543
4427 4424
6748 6739
6792 6781
9494 9475
10239 10216
12656 12623
14209 14168
16150 16107
16226 16173
18673 18606
20195 20116
21414 21338
23611 23510
24559 24708
27203 27084
29006 28859
29180 29053
29309 29151

Take a look at the attached PDF file for the shape of all the modes, including all that are not visible. We see all the modes we expect to be able to excite with the central suspension of the disk.

The roughing pump is making a lot of non stationary low frequency noise. I turned it off, and the pressure stayed constant at 1.2e-6 Torr over about 1.5 hours. Here's the difference in the QPD spectrum:

It turned out that I have enough excitation authority to knock the disk out of the right place. So I had to vent to recover the situation. I'll open the chamber tomorrow and see what happened.

Here's a first bird eye look at the ring downs. We see beating of the two almost denegerate modes in some cases. Fits will follow, using the procedure I used for the LMA measurements.

Attachment 1: modes.pdf
modes.pdf modes.pdf modes.pdf modes.pdf modes.pdf modes.pdf modes.pdf modes.pdf
  91   Fri Aug 19 19:05:01 2016 GabrieleOpticsDaily ProgressDisk installed into the chamber

I assembled the disk suspension sytem and installed into the chamber. Although I don't have the magnets and coils, I installed the movable retaining disk, and used it to center the disk.

I first aligned the input laser using the reflection off the black glass, which turns out to be quite bright and very well visible. Tomorrow I'm going to measure how much power we have in the black glass.

The reflection from the disk is slighlty separated from the reflection from the black glass, so I can block it using an iris.

At 6:50pm I closed the chamber and started the roughing pump. At 7:05pm pressure was below 1 Tor so I started the turbo pump. When leaving pressure is about 1.6e-5 Tor.

  90   Fri Aug 19 13:46:55 2016 Alena, Calum, Jon, Liz, Steve, GabrieleFacilityDaily ProgressClean room and vacuum complete

This morning we installed the clean room curtains and washed them. It turns out that the air filters are supposed to be powered at 277V (?) instead of 115V. So right now the flux is quite low. We are looking into the problem: either replace them with 115V modules or install a small transformer.

We also installed the vacuum chamber on the table and connected all the pumps and gauges. There are no leaks and we could pump down easily the empty chamber. We left for lunch when the pressure was at a few 1e-6 Tor and still going down.

  89   Thu Aug 18 18:15:18 2016 Gabriele, AlenaCleanDaily ProgressClean room construction progress

Ceiling, back and side panels are installed. The air filters have been cabled and connected to the power supply.

  88   Wed Aug 17 16:41:05 2016 Gabriele, AlenaCleanDaily ProgressClean room construction progress

The clean room frame is built and secured to the floor and wall. Panels are being installed on the ceiling and back. Also, the optical table has been leveled.

  87   Tue Aug 16 11:57:03 2016 GabrieleElectronicsConfigurationFixed cymac3 network and time issue

The cymac3 internal clock was off by about 10 seconds. When I tried to start the NTP service, I found out that the cymac3 couldn't reach any external server. It turned out that the gateway in /etc/network/interfaces was set to the wrong address. I fixed it and rebooted. Now NTP is working and the time is correct.

This fixed a small issue with diaggui, which always complained about a data receiving error when starting a measurement (although after the complian the measurement could continue)

  86   Tue Aug 16 11:53:30 2016 Alena, GabrieleCleanDaily ProgressClean room progress

Some progress on the cleam room: bar fixed to the wall, some more structure built, filters in place. We had to (literally) work around a corner of the low ceiling that we haven't noticed before. More contruction will follow tomorrow. We also had to order some additional parts (more extrusions, brackets, screws, etc...) 

  85   Sun Aug 14 18:26:19 2016 GabrieleGeneralGeneralPlans for the automatic mode search

Here's how I imagine the mode search to proceed:

  1. the user select the list of nominal frequencies and the frequency range for the search
  2. the automation creates an excitation which is a sum of band-limited white noise, centered around each of the modes, with the desired bandwidth. The excitation is normalized to match a reasonable fraction of the DAC range
  3. the excitation is injected for a given amount of time and each band is searched for the mode. In each band, if a mode is found, the frequency is estimated together with the maximum SNR obtained and the results are saved
  4. if some modes are not found, the automation will try to improve the noise, by scaling all band amplitudes in such a way to have the same resulting SNR for the modes. In this way we should get more dynamical range and could increase the excitation for the modes that were not found
  5. if some modes are still unidentified, then the user will be asked to choose between a few options:
    1. skip the mode
    2. try to excite it with an excitation focused on that band only
    3. increase the bandwidth for the search
    4. maintain the same bandwidth but move the band to adjacent frequencies
  6. At the end all identified modes will be stored in memory and saved to a file defined by the user
  84   Sun Aug 14 18:13:52 2016 GabrieleElectronicsDaily ProgressProgress on the user interface

Some progress on writing the user interface:

Now the user can open a file that defines the nominal mode frequencies (from COMSOL simulations) and select which modes to search for:

The plan is that the automation will then inject band limited noise around each nominal frequency, to excite the mode, and then find the exact frequency. The user can set some additional parameters like

  • the duration of the excitation (I plan to use this as a maximum: if the mode rings up with large enough SNR after a shorter time, the injection will stop earlier)
  • the amplitude of the excitation (units to be determined)
  • the bandwitdh, otr in other words how many Hz the noise will be spread around the nominla frequency to search for the mode.
  • the minimum SNR to detect a peak (here the automation will get a spectrum before any excitation, to be used as the baseline to whiten all further spectra)
  83   Sun Aug 14 11:40:29 2016 GabrieleElectronicsDaily ProgressSome python code

Today I wrote some auxiliary functions that will be useful for the measurement system:

  • noise.py: set of functions to generate band-limited noise (using inverse FFT) and multiple band noise. Using the awg python interface I can also start and stop the noise injection. Some examples of the result:

     
  • readdata.py: read data online, compute a spectrum, and additionally a function that find peaks in a whitened spectrum. All peaks above a minimum SNR are returned: the central frequency is computed by an average of the bin SNR:

     
  • diskmodel.py: reads a list of mode frequencies from txt files. In each file, the first row is the disk diameter in mm, the second is the disk thickness in mm, all other rows are the modes are computed by COMSOL. I also produced a whole bunch of such files, with diameters within 75 +- 0.1 mm and thickness within 1 +- 0.1 mm
  82   Sat Aug 13 14:05:02 2016 GabrieleGeneralDaily ProgressFirst work on the measurement user interface

Today I started programming part of the user interface that will be used to perform the measurements. Not much implemented so far, but you can get an idea of the look:

Buttons on the left sidebar will allow the user to perform some basic tasks. The main panel has a plot (which will show spectra or ring down measurements) and a log section.

Attachment 1: Screenshot-C.Ri.Me..png
Screenshot-C.Ri.Me..png
ELOG V3.1.3-